PLEASE HELPP! A large stock traded on Borsa Istanbul can realize a positive or a negative return independently on a given day.​ (Assume that every day there are many trades​executed, so that the return is never​ zero.) A positive return means that the stock ended the day with a closing price​ increase, while a negative return means that the stock ended the day with a closing price​ decrease, compared to the previous trading day. Suppose that closing price increases and closing price decreases are​equally-likely events. What is the chance that the Borsa Istanbul stock will have an increase in its closing price on seven consecutive trading​ days ?

Note : Use Excel or the cumulative probabilities table to compute the probability.

The chance that the Borsa Istanbul stock would have an increase in its closing price on seven consecutive trading days is ... ?  ​(Round to four decimal​ places.)
Answer #1

S = X1 + X2+ X3+..+X7

Xi = 1 with p = 0.5

= -1 with p = 0.5

P(S > 0)

when there will be greater than or equal to 4 success (Xi = 1}

S follow binomial distribution with n = 7 and p = 0.5

P(S >= 4)

= 1- P(S <= 3)

= 1- binom.dist(3,7,0.5,1)

= 0.5